# Gravitational signatures of a nonlinear electrodynamics in $f(R,T)$ gravity

### v1

@@ Abstract
In this work, we investigate a nonlinear electrodynamics model within the framework of $f(R,T)$ gravity. We begin by outlining the general features of the theory and analyzing the event horizon under conditions ensuring its real and positive definiteness. We then examine light trajectories, focusing on critical orbits, shadow radii, and geodesics of massless particles. The parameters $\alpha$ and $\beta$, associated with the nonlinear extension of the Reissner–Nordström spacetime, are constrained using observational data from the Event Horizon Telescope (EHT). Subsequently, we analyze the thermodynamic properties of the system, including Hawking temperature, entropy, and heat capacity. Quasinormal modes are computed for scalar, vector, tensor, and spinorial perturbations, with the corresponding time–domain profiles explored as well. Gravitational lensing is then studied in both weak and strong deflection limits, along with the stability of photon spheres. Finally, we examine additional topological aspects, including topological thermodynamics and the topological photon sphere.
@@ I Introduction
Extensions of General Relativity (GR) have garnered considerable attention in recent years, primarily due to their potential to resolve key observational inconsistencies within the $\Lambda$CDM framework. These include the persistent $H_{0}$ tension Hu:2023jqc , the $S_{8}$ discrepancy Joseph:2022jsf , and emerging evidence from DESI indicating a possible variation in the cosmological constant DESI:2025zgx . In addition to observational motivations, several theoretical challenges—such as the occurrence of singularities Senovilla:1998oua , the unresolved nature of dark energy Frusciante:2019xia , and the underlying mechanism driving cosmic inflation Vazquez:2018qdg —have further spurred interest in modifications to GR. Beyond modifications arising from additional terms in the Einstein–Hilbert action—motivated by various theoretical arguments—recent studies have explored the possibility of relaxing the coupling between gravity and matter as a means to introduce novel effects that may help resolve the aforementioned tensions in GR BarrosoVarela:2024htf . This approach, known as non–minimal coupling, can be implemented in various forms (for a comprehensive review, see Velten:2021xxw ).
@@
Extreme environments offer a valuable testing ground for examining the boundaries of fundamental physical theories. In particular, nonlinear extensions of Maxwell electrodynamics have garnered significant interest for their potential to resolve singularities in both cosmological and astrophysical settings PhysRevD.65.063501 ; Sorokin:2021tge . Among various astrophysical phenomena, black hole observables have become especially promising tools for probing fundamental physics—an interest amplified by the landmark detections from the LIGO and VIRGO collaborations LIGOScientific:2016aoc , as well as high–resolution imaging provided by the Event Horizon Telescope (EHT) collaboration EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp .
@@
Gravitational wave astronomy has emerged as a powerful tool for probing a wide range of phenomena, including gravitational lensing effects heidari2023gravitational ; 020 ; araujo2024gravitational ; 019 ; aa2024implications ; araujo2023analysis . Traditionally, research in this area has focused on light deflection in weak gravitational fields, particularly within the context of the Schwarzschild metric and more general static, spherically symmetric spacetimes 021 . However, the strong–field regime near compact objects such as black holes is anticipated to exhibit pronounced deviations from classical predictions, potentially enabling critical tests of both general relativity and nonlinear electrodynamics 022 .
@@
Black hole shadows serve as powerful observational features for investigating the behavior of gravity in the strong-field regime. These dark regions, set against the backdrop of luminous emission from accretion of matter, result from extreme light deflection near the event horizon and carry information about the geometry of the surrounding spacetime. Theoretical studies of this phenomenon date back to Bardeen in the 1970s Cunningham , with further developments by Falcke, Melia, and Agol Falcke:1999pj , who proposed that the shadow of $SgrA^{*}$ could be detectable via very long baseline interferometry operating at submillimeter wavelengths. This prediction was confirmed in 2019, when the Event Horizon Telescope (EHT) collaboration released the first image of a black hole in the center of the galaxy $M87$, followed by the imaging of $SgrA^{*}$. These results marked a turning point in gravitational physics, enabling detailed comparisons between general relativity and alternative theories in highly curved spacetimes Vagnozzi:2019apd ; Bambi:2019tjh ; Allahyari:2019jqz ; Kumar:2020hgm ; Afrin:2021imp ; Khodadi:2021gbc ; Afrin:2021wlj ; Khodadi:2022pqh ; Fu:2021fxn ; Afrin:2022ztr ; Afrin:2023uzo ; Ghosh:2022kit ; Afrin:2024khy ; Khodadi:2024ubi .
@@
Moreover, small disturbances in the region surrounding a black hole lead to characteristic oscillatory responses, which dominate the gravitational wave signal during the phase known as the ringdown Konoplya:2019hlu ; Konoplya:2013rxa ; Kokkotas:2010zd ; Konoplya:2007zx ; Konoplya:2011qq . These oscillations occur at complex frequencies referred to as quasinormal modes, which depend solely on the black hole’s intrinsic parameters. The real part of each mode determines the oscillation frequency, while the imaginary component controls the rate at which the amplitude decays over time. Because quasinormal spectra are directly linked to the black hole’s mass, charge, and angular momentum, they offer a precise characterization of the object and also establish a connection with its shadow properties Jusufi:2020dhz and the greybody factors Konoplya:2024vuj ; Konoplya:2024lir .
@@ II The black hole solution and the general features
Notice that these conditions must be fulfilled simultaneously. Fig. 1 provides a qualitative illustration, while Table 1 presents the corresponding quantitative values of the event horizon $r_{h}$. Both indicate that increasing the magnetic charge $Q$—with fixed parameters $\alpha=\beta=-0.01$—leads to a reduction in the event horizon radius. Additionally, for fixed $Q$, a decrease in the coupling parameters $\alpha=\beta$ results in a smaller value of $r_{h}$.
@@ III.2 Photon spheres and shadows
$-\mathrm{A}(r){{\dot{t}}^{2}}+\mathrm{B}(r){{\dot{r}}^{2}}+\mathrm{D}(r){{\dot{\varphi}}^{2}}=0.$ (15)
@@
[…]
@@ III.3 Bounds to shadows based on EHT data
$5.564.21<\frac{\mathcal{R}}{M}<$ 5.56 . (25)
Fig. 4 presents the graphical constraints on the charge $Q$ for various fixed values of the parameters $\alpha$ and $\beta$. It is worth noting that the reverse procedure—varying $\alpha$ and $\beta$ for different fixed values of $Q$—was not carried out in this analysis, as the resulting curves become visually indistinguishable, making interpretation challenging in that configuration. In addition, this is why we also have considered “huge” values for $\alpha$ and $\beta$. To complement this analysis, Table 4 is included to offer a detailed quantitative assessment of the bounds derived in this work.
@@ IV.1 Hawking temperature
[…]
@@ V.1 Scalar perturbations
$-\frac{\partial^{2}\tilde{\Psi}}{\partial t^{2}}+\frac{\partial^{2}\tilde{\Psi}}{\partial r^{*2}}+V_{\text{S}}(r)\tilde{\Psi}=0.$ (38)
@@
Tabs. 5 and 6 compile the complex quasinormal frequencies for multipole numbers $\ell=1$ and $\ell=2$, considering various parameter choices. When the charge is fixed at $Q=0.99$, reducing $\alpha=\beta$ results in alternating behavior between weaker and stronger damping across all modes $\omega$ for both $\ell=1$ and $\ell=2$. A similar pattern emerges as $Q$ increases with $\alpha=\beta$ held constant at $-0.01$, producing an “irregular” sequence of less and more damped frequencies, with the exception of $\omega_{2}$, which does not follow this trend. In other words, this specific mode exhibits reduced damping for the latter configuration (when $Q$ runs for fixed values of $\alpha$ and $\beta$).
@@ V.3 Tensor perturbations
Tabs. 9 and 10 report the quasinormal frequencies for tensor perturbations with $\ell=1$ and $\ell=2$, covering a range of values for $Q$, $\alpha$, and $\beta$. As with earlier cases, the damping rates alternate depending on the chosen parameters, exhibiting either stronger or weaker attenuation. A distinct behavior emerges here compared to the scalar and vector sectors: for the mode $\omega_{2}$, keeping $\alpha$ and $\beta$ fixed while varying $Q$ may lead to instabilities in the computed frequencies—at least within the parameter explored here. It is also important to point out that, specifically for the case $\ell=2$, the third–order WKB approximation was adopted, in contrast to the sixth–order scheme utilized earlier for scalar and vector perturbations.
@@ VII Time–domain solution: bosonic case
To construct the numerical grid, points are designated within the null coordinate plane as follows: the base point is $\tilde{S}=(\tilde{u},\tilde{v})$, while neighboring positions are labeled $\tilde{W}=(\tilde{u}+\tilde{h},\tilde{v})$, $\tilde{E}=(\tilde{u},\tilde{v}+\tilde{h})$, and $\tilde{N}=(\tilde{u}+\tilde{h},\tilde{v}+\tilde{h})$, where $\tilde{h}$ denotes the step size used for discretization. The computational domain is initialized along the characteristic lines $\tilde{u}=\tilde{u}_{0}$ and $\tilde{v}=\tilde{v}_{0}$, which serve as the foundation for evolving the system forward. As a starting profile, a Gaussian function centered at $\tilde{v}=\tilde{v}_{c}$ with width parameter $\sigma$ is imposed along the initial slice $\tilde{u}=\tilde{u}_{0}$, specifying the initial wave configuration for the evolution algorithm
@@ VII.1 Scalar perturbations
To examine the attenuation more precisely, Fig. 14 plots the logarithm of the absolute value, $\ln|\tilde{\psi}|$, for the same set of charges and multipole indices. The damping pattern is preserved, and a transition to late–time power–law decay becomes evident—consistent with the expected tail behavior following the quasinormal phase.
@@
Figure 14: Logarithmic profile of the scalar field amplitude, $\ln|\tilde{\psi}|$, as a function of time $t$ for fixed $\alpha=\beta=-0.01$ and varying charge values $Q=0.6$, $0.7$, $0.8$, and $0.9$. Results are shown for angular modes $\ell=0$ (top left), $\ell=1$ (top right), and $\ell=2$ (bottom).
Figure 15: The figure presents the evolution of the scalar field in a ln–ln scale, where $\ln|\tilde{\psi}|$ is plotted against $\ln|t|$ for several values of the charge parameter $Q$, while the parameters $\alpha=\beta$ remain fixed at $-0.01$. The charge values examined are $Q=0.6$, $0.7$, $0.8$, and $0.9$. Each subplot corresponds to a distinct multipole number: $\ell=0$ in the upper–left frame, $\ell=1$ in the upper–right, and $\ell=2$ in the lower panel.
@@ VII.2 Vector perturbations
A comprehensive investigation of the temporal evolution of vector field disturbances is conducted by examining how the signal $\tilde{\psi}$ behaves under variations of the electric charge parameter $Q$, while maintaining fixed coupling values $\alpha=\beta=-0.01$. As illustrated in Fig. 16, the analysis spans four distinct values of $Q$—namely $0.6$, $0.7$, $0.8$, and $0.9$—and considers angular momentum numbers , $\ell=1$, and $2$, depicted in the left, and right panels, respectively. In all configurations, the waveform undergoes damped oscillations, gradually losing amplitude as time progresses.
To highlight the attenuation pattern more effectively, Fig. 17 presents a logarithmic plot where the vertical axis displays $\ln|\tilde{\psi}|$ against time $t$. The decay curves demonstrate that the initial exponential suppression of the signal transitions into a slower falloff at later stages. This shift signals the emergence of a power–law tail—a late–time signature often arising from scattering effects in curved backgrounds.
@@
Figure 17: Logarithmic profile $\ln|\tilde{\psi}|$ for vector–type perturbations plotted against time $t$, considering $Q=0.6$, $0.7$, $0.8$, and $0.9$, with coupling parameters set to $\alpha=\beta=-0.01$. The left and right panels display the results for $\ell=1$ and $\ell=2$, respectively.
Figure 18: $\ln$$\ln$ plots of $\ln|\tilde{\psi}|$ versus $\ln|t|$ for vector perturbations, shown for charge values $Q=0.6$, $0.7$, $0.8$, and $0.9$, with the parameters $\alpha=\beta$ kept constant at $-0.01$. The left panel corresponds to $\ell=1$, while the right panel presents results for $\ell=2$.
@@ VII.3 Tensor perturbations
This part addresses the temporal response of tensorial fluctuations under varying electric charge. The waveform $\tilde{\psi}$ is exhibited as it evolves with time $t$, as shown in Fig. 19, where the influence of different $Q$ values—specifically $0.6$, $0.7$, $0.8$, and $0.9$—is explored while the deformation parameters remain fixed at $\alpha=\beta=-0.01$. The plots correspond to $\ell=1$ (left panel) and $\ell=2$ (right panel). Furthermore, in order to show another detailed view of this decay process, we show Fig. 20, which it is displayed $\ln|\tilde{\psi}|$ versus time $t$. To emphasize the asymptotic regime, Fig. 21 employs a double–logarithmic scale, plotting $\ln|\tilde{\psi}|$ against $\ln|t|$ for the same combinations of $Q$ and angular index.
@@
Figure 20: Time evolution of tensor–mode perturbations depicted through the logarithmic profile $\ln|\tilde{\psi}|$ as a function of $t$, evaluated for charge values $Q=0.6$, $0.7$, $0.8$, and $0.9$, with fixed parameters $\alpha=\beta=-0.01$. The plots correspond to angular indices $\ell=1$ (left) and $\ell=2$ (right).
Figure 21: Late–time behavior of tensorial perturbations displayed through a $\ln|\tilde{\psi}|$ versus $\ln|t|$ plot, where the coupling constants are fixed at $\alpha=\beta-0.01$ and the electric charge parameter varies over $Q=0.6$, $0.7$, $0.8$, and $0.9$. The left and right panels correspond to angular momentum numbers $\ell=1$ and $\ell=2$.

### v2

@@ Abstract
In this work, we investigate a nonlinear electrodynamics model in the context of $f(R,T)$ gravity. We begin by outlining the general features of the theory and analyzing the event horizon under conditions ensuring its real and positive definiteness. We then examine light trajectories, focusing on critical orbits, shadow radii, and geodesics of massless particles. The parameters $\alpha$ and $\beta$, associated with the nonlinear extension of the Reissner–Nordström spacetime, are constrained using observational data from the Event Horizon Telescope (EHT). Subsequently, we analyze the thermal aspects of the system, including Hawking temperature, entropy, and heat capacity. Quasinormal modes are computed for scalar, vector, tensor, and spinorial perturbations, with the corresponding time–domain profiles explored as well. Gravitational lensing is then studied in both weak and strong deflection limits, along with the stability of photon spheres. Finally, we examine additional topological aspects, including topological thermodynamics and the topological photon sphere.
@@ I Introduction
[…]
@@
Extreme environments offer a valuable testing ground for examining the boundaries of fundamental physical theories. In particular, nonlinear extensions of Maxwell electrodynamics have garnered significant interest for their potential to resolve singularities in both cosmological and astrophysical settings PhysRevD.65.063501 ; Sorokin:2021tge . In fact, some physical processes indicate that we should modify Maxwell’s electromagnetism. For instance, vacuum polarization induces small deviations from QED, such as birefringence Battesti:2012hf , which could be phenomenologically described as effects of nonlinear electrodynamics Sorokin:2021tge ; Boillat:1970gw ; Bialynicka-Birula:1970nlh . From a theoretical perspective, nonlinear modifications of Maxwell electrodynamics arise in quantum gravity theories, such as ModMax theory, string theory and M–theory Gibbons:2001gy ; Bandos:2020hgy ; heidari2025gravitational ; 2024arXiv241102907E ; panah2025some , and have been investigated in a myriad of scenarios, from cosmology and astrophysics to condensed matter physics Ayon-Beato:2000mjt ; Pereira:2018mnn ; Gaete:2014nda ; Novello:2006ng ; Pan:2011vi ; Neves:2023uxi . Experimental efforts have been undertaken to search for traces of these effects, including Pb–Pb collisions at the LHC ATLAS:2017fur and investigations of light–by–light scattering and photon splitting at the ROKK-1M facility Akhmadaliev:2001ik .
@@
Among various astrophysical phenomena, black hole observables have become especially promising tools for probing fundamental physics—an interest amplified by the landmark detections from the LIGO and VIRGO collaborations LIGOScientific:2016aoc , as well as high–resolution imaging provided by the Event Horizon Telescope (EHT) collaboration EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp .
@@
Gravitational wave astronomy has emerged as a powerful tool for probing a wide range of phenomena, including gravitational lensing effects heidari2023gravitational ; 020 ; araujo2024gravitational ; 019 ; aa2024implications ; araujo2023analysis . Traditionally, research in this area has focused on light deflection in weak gravitational fields, particularly within the context of the Schwarzschild metric and more general static, spherically symmetric spacetimes 021 . However, the strong–field regime near compact objects such as black holes is anticipated to exhibit pronounced deviations from classical predictions, potentially enabling critical tests of both general relativity and nonlinear electrodynamics 022 ; mohan2025strong .
@@
Black hole shadows serve as powerful observational features for investigating the behavior of gravity in the strong-field regime. These dark regions, set against the backdrop of luminous emission from accretion of matter, result from extreme light deflection near the event horizon and carry information about the geometry of the surrounding spacetime. Theoretical studies of this phenomenon date back to Bardeen in the 1970s Cunningham , with further developments by Falcke, Melia, and Agol Falcke:1999pj , who proposed that the shadow of $SgrA^{*}$ could be detectable via very long baseline interferometry operating at submillimeter wavelengths. This prediction was confirmed in 2019, when the Event Horizon Telescope (EHT) collaboration released the first image of a black hole in the center of the galaxy $M87$, followed by the imaging of $SgrA^{*}$. These results marked a turning point in gravitational physics, enabling detailed comparisons between general relativity and alternative theories in highly curved spacetimes Vagnozzi:2019apd ; Bambi:2019tjh ; Allahyari:2019jqz ; Kumar:2020hgm ; Afrin:2021imp ; Khodadi:2021gbc ; Afrin:2021wlj ; Khodadi:2022pqh ; Fu:2021fxn ; Afrin:2022ztr ; Afrin:2023uzo ; Ghosh:2022kit ; Afrin:2024khy ; Khodadi:2024ubi ; Liu:2024lve ; Nojiri:2024qgx ; Nojiri:2024nlx ; Nojiri:2024txy ; karmakar2023thermodynamics .
@@
Moreover, small disturbances in the region surrounding a black hole lead to characteristic oscillatory responses, which dominate the gravitational wave signal during the phase known as the ringdown Konoplya:2019hlu ; Konoplya:2013rxa ; Kokkotas:2010zd ; karmakar2024quasinormal ; gogoi2023quasinormal ; karmakar2022quasinormal ; Konoplya:2007zx ; gogoi2024quasinormal ; Konoplya:2011qq . These oscillations occur at complex frequencies referred to as quasinormal modes, which depend solely on the black hole’s intrinsic parameters. The real part of each mode determines the oscillation frequency, while the imaginary component controls the rate at which the amplitude decays over time. Because quasinormal spectra are directly linked to the black hole’s mass, charge, and angular momentum, they offer a precise characterization of the object and also establish a connection with its shadow properties Jusufi:2020dhz and the greybody factors Konoplya:2024vuj ; Konoplya:2024lir . Recently, it has been claimed that quasinormal modes have been detected, but this is still a matter of debate where the control of uncertainties play a significant role Franchini:2023eda . We expect that future runs of the LIGO/Virgo/KAGRA collaboration can help to clarify this issue.
@@ II The black hole solution and the general features
Notice that these conditions must be fulfilled simultaneously. Fig. 1 provides a qualitative illustration, while Table 1 presents the corresponding quantitative values of the event horizon $r_{h}$. Both indicate that increasing the magnetic charge QQ—with fixed parameters $\alpha=\beta=-0.01$—leads to a reduction in the event horizon radius. Additionally, for fixed $Q$, a decrease in the coupling parameters $\alpha=\beta$ results in a smaller value of $r_{h}$.
@@ III.1 Geodesics
On the other hand, Fig. 3 illustrates how the parameters $\alpha$ and $\beta$ affect the trajectory of light. Overall, for $Q=0.5$, we observe that decreasing these parameters leads to a weaker deflection, according to the numerical conditions adopted here.
@@
Figure 3: The numerically obtained geodesics are shown for $\alpha=\beta=-0.01$, 0.1-0.1, 0.2-0.2, and 0.3-0.3, with the charge fixed at $Q=0.5$. The large colored disk represents the event horizon, and the dot–dashed lines mark the photon sphere corresponding to the chosen parameter values.
@@ III.2 Photon spheres and shadows
$\mathrm{A}(r){{\dot{t}}^{2}}+\mathrm{B}(r){{\dot{r}}^{2}}+\mathrm{D}(r){{\dot{\varphi}}^{2}}=0.$ (15)
@@
$Q$ $\alpha=\beta$ $\mathcal{R}$ $Q$ $\alpha=\beta$ $\mathcal{R}$
@@ III.3 Bounds to shadows based on EHT data
$4.21<\frac{\mathcal{R}}{M}<5.56.$ (25)
Fig. 5 presents the graphical constraints on the charge $Q$ for various fixed values of the parameters $\alpha$ and $\beta$. It is worth noting that the reverse procedure—varying $\alpha$ and $\beta$ for different fixed values of QQ—was not carried out in this analysis, as the resulting curves become visually indistinguishable, making interpretation challenging in that configuration. In addition, this is why we also have considered “huge” values for $\alpha$ and $\beta$. To complement this analysis, Table 4 is included to offer a detailed quantitative assessment of the bounds derived in this work.
@@ IV.1 Hawking temperature
$\begin{split}M_{rem}&=\frac{\sqrt{\frac{5\sqrt[3]{3}Q^{4}+\sqrt[3]{5}\left(9\alpha(1-2\beta)Q^{4}+\sqrt{3}\sqrt{Q^{8}\left(27\alpha^{2}(1-2\beta)^{2}-25Q^{4}\right)}\right)^{2/3}}{\sqrt[3]{9\alpha(1-2\beta)Q^{4}+\sqrt{3}\sqrt{Q^{8}\left(27\alpha^{2}(1-2\beta)^{2}-25Q^{4}\right)}}}}}{2\sqrt[3]{15}}\\ &\approx\,\frac{Q}{2}+\frac{\alpha(1-2\beta)}{20Q}.\end{split}$ (31)
@@ V.1 Scalar perturbations
$\frac{\partial^{2}\tilde{\Psi}}{\partial t^{2}}+\frac{\partial^{2}\tilde{\Psi}}{\partial r^{*2}}+V_{\text{S}}(r)\tilde{\Psi}=0.$ (38)
@@
Tabs. 5 and 6 compile the complex quasinormal frequencies for multipole numbers $\ell=1$ and $\ell=2$, considering various parameter choices. When the charge is fixed at $Q=0.99$, reducing $\alpha=\beta$ results in alternating behavior between weaker and stronger damping across all modes $\omega$ for both $\ell=1$ and $\ell=2$. A similar pattern emerges as $Q$ increases with $\alpha=\beta$ held constant at 0.01-0.01, producing an “irregular” sequence of less and more damped frequencies, with the exception of $\omega_{2}$, which does not follow this trend. In other words, this specific mode exhibits reduced damping for the latter configuration (when $Q$ runs for fixed values of $\alpha$ and $\beta$).
@@ V.3 Tensor perturbations
Tabs. 9 and 10 report the quasinormal frequencies for tensor perturbations with $\ell=1$ and $\ell=2$, covering a range of values for $Q$, $\alpha$, and $\beta$. As with earlier cases, the damping rates alternate depending on the chosen parameters, exhibiting either stronger or weaker attenuation. A distinct behavior emerges here compared to the scalar and vector sectors: for the mode $\omega_{2}$, keeping $\alpha$ and $\beta$ fixed while varying $Q$ may lead to instabilities in the computed frequencies—at least within the parameter explored here. It is also important to point out that, specifically for the case $\ell=2$, the third–order WKB approximation was adopted, in contrast to the sixth–order scheme utilized earlier for scalar and vector perturbations. Recent claims on the detection of QNMs beyond the $\ell=m=2$ have been a matter of recent discussion and we expect that this window could be used to constrain alternative gravity proposals in the future Franchini:2023eda .
@@ VII Time–domain solution: bosonic case
[…]
@@ VII.1 Scalar perturbations
[…]
@@
[…]
@@ VII.2 Vector perturbations
A comprehensive investigation of the temporal evolution of vector field disturbances is conducted by examining how the signal $\tilde{\psi}$ behaves under variations of the electric charge parameter $Q$, while maintaining fixed coupling values $\alpha=\beta=-0.01$. As illustrated in Fig. 17, the analysis spans four distinct values of QQ—namely $0.6$, $0.7$, $0.8$, and 0.90.9—and considers angular momentum numbers , $\ell=1$, and $2$, depicted in the left, and right panels, respectively. In all configurations, the waveform undergoes damped oscillations, gradually losing amplitude as time progresses.
[…]
@@
[…]
@@ VII.3 Tensor perturbations
[…]
@@
[…]